Presented below is a partially completed Statement of Activities for a homeless shelter. Complete the Statement of Activitie how would i solve the contribution revenues

am i missing anythin?
Answer #1

The completed statement of activities is provided as follows:

CENTERVILLE AREA HOMELESS SHELTER
Statement of Activities
For the Year Ended December 31, 2020
Net Assets without
Donor Restrictions
Net Assets with
Donor Restrictions
Total
Revenues
Contribution Revenues 1,528,800 53,710 1,582,510
Net Assets Released from Restrictions 27,000 -27,000 0
Total Revenues 1,555,800 26,710 1,582,510
Expenses
Temporary Shelter Program 1,028,100 0 1,028,100
Self-Sufficiency Program 365,000 0 365,000
Fund Raising 7,750 0 7,750
Administration 70,950 0 70,950
Total Expenses 1,471,800 0 1,471,800
Increase in Net Assets 84,000 26,710 110,710
Net Assets January 1 21,000 14,350 35,350
Net Assets December 31 $105,000 $41,060 $146,060

_____

Notes:

The calculations are given as below:

Net Assets without Donor Restrictions

Total Revenues (Net Assets without Donor Restrictions) = Increase in Net Assets (Net Assets without Donor Restrictions) + Total Expenses (Net Assets without Donor Restrictions) = 84,000 + 1,471,800 = $1,555,800

Contribution Revenues (Net Assets without Donor Restrictions) = Total Revenues (Net Assets without Donor Restrictions) - Net Assets Released from Restrictions (Net Assets without Donor Restrictions) = 1,555,800 - 27,000 = $1,528,800

____

Net Assets with Donor Restrictions

Total Revenues (Net Assets with Donor Restrictions) = Increase in Net Assets (Net Assets with Donor Restrictions) + Total Expenses (Net Assets with Donor Restrictions) = 26,710 + 0 = $26,710

Contribution Revenues (Net Assets with Donor Restrictions) = Total Revenues (Net Assets with Donor Restrictions) - Net Assets Released from Restrictions (Net Assets with Donor Restrictions) = 26,710 - (-27,000) = $53,710